# Strand, London WC2R 2LS, UK

### v1

@@ 1 Introduction
This intuition can be made precise in various top-down constructions of AdS/CFT from string theory, the prototypical example being the duality [^1] At large-$N$ and large ’t Hooft coupling $\lambda$. between a type IIB supergravity AdS black brane and $\mathcal{N}=4$ super Yang-Mills with $SU(N)$ gauge group at finite temperature Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj ; Witten:1998zw . In this example, one finds $N^{2}\propto\ell_{\textrm{\tiny AdS}}^{d-1}/G_{N}$.
From the bulk point of view, the thermodynamic behavior of black holes can be analyzed with the Euclidean gravitational path integral via the standard Gibbons-Hawking prescription Gibbons:1976ue . This prescription can also accommodate situations where one is interested in (quasi-local) thermodynamics of finite spacetime regions, by for instance placing the black hole inside a finite boundary as in the classic work of York York:1986it . In that particular case, spherical Dirichlet boundary conditions were imposed at the boundary.
Surprisingly enough, however, Dirichlet boundary conditions, where one fixes the profile of the induced metric along the boundary $\partial\mathcal{M}$, do not generically yield a well-posed (Initial) Boundary Value Problem (IBVP) for general relativity (GR) in (Lorentzian) Euclidean signature anderson2008boundary ; anderson2010extension ; Anninos:2022ujl . Furthermore, in Euclidean signature, they also fail to be elliptic, meaning that quantum perturbation theory around a classical solution is ill-defined Avramidi:1997sh ; Witten:2018lgb . [^2] Note, however, that certain classes of boundary geometries exhibit better behavior under Dirichlet boundary conditions, provided the Brown–York stress tensor defines a Lorentzian metric on the boundary of the same signature as the bulk An:2025gvr . An analogue statement holds in Euclidean signature anderson2010extension ; Witten:2018lgb . These include e.g., spherical boundaries in de Sitter space Coleman:2021nor ; Silverstein:2024xnr .
An interesting choice of boundary conditions are the conformal boundary conditions (CBC) York:1972sj , whereby the conformal structure of the boundary $h_{mn}]$ along with the trace of the extrinsic curvature $\mathcal{K}$ are held fixed,
@@
Conformal boundary conditions yield a well-posed and elliptic Euclidean boundary problem anderson2008boundary ; anderson2010extension . The Lorentzian problem is more intricate, but has better existence and uniqueness properties than the Dirichlet case. See, e.g., An:2020nfw ; An:2021fcq ; An:2025rlw ; An:2025gvr ; An:2025cbs ; Liu:2025xij for recent developments.
The thermodynamic properties of black holes in Einstein gravity subject to CBC have been investigated with $d=2$ and 3 in Anninos:2023epi ; Anninos:2024wpy ; Galante:2025tnt and the analysis was then generalized to arbitrary dimension $d$ in Banihashemi:2024yye ; Banihashemi:2025qqi . When discussing Euclidean thermodynamics subject to CBC, the canonical thermal ensemble is defined by a conformally invariant temperature $\tilde{\beta}$ and the trace of the extrinsic curvature $\mathcal{K}$ and hence the Euclidean problem gets modified, compared to its Dirichlet counterpart. One of the main outcomes of Anninos:2023epi ; Anninos:2024wpy ; Galante:2025tnt ; Banihashemi:2024yye ; Banihashemi:2025qqi was that, in the small-$\tilde{\beta}$ limit, the conformal entropy has a structure that coincides with the expectations for the thermal effective action of a local CFT in one lower dimension – in contrast with Dirichlet boundary conditions. More specifically, the leading contribution to the black hole conformal entropy in the high-temperature expansion was found to be
@@
where $\Lambda=\pm\tfrac{d(d-1)}{2\ell_{\textrm{\tiny(A)dS}}^{2}}$ and $\Sigma_{k,d-1}$ is the volume of the (d1)(d-1)-dimensional maximally symmetric black hole horizon. Note that in $d=2$, and for black branes, these formulas are exact. This observation, drawing inspiration from the traditional AdS/CFT examples (1), could possibly provide hints towards a version of holography for finite portions of spacetime, even beyond $\Lambda<0$. A tree-level proposal for the dual field theoretic description of these timelike boundaries in AdS3, that reproduces eq. (5), was recently proposed in Allameh:2025gsa .
Given the above, a well-motivated question is to study the universality of the extensive behavior of the conformal entropy (4), when higher curvature contributions to the gravitational effective action are taken into account. In this work, we take a step towards addressing this question by considering Einstein-Maxwell-Gauss-Bonnet (EMGB) gravity. Further, we provide arguments regarding the robustness of our conclusions against further higher curvature terms of Lanczos-Lovelock (LL) type Lanczos:1938sf ; Lovelock:1971yv .
@@
Studying the thermodynamics of EMGB black holes with spherical horizons subject to the ECBC, we find that the extensive form of the conformal entropy in the high-temperature limit (4) persists, with (5) generalizing to
$\mathcal{N}_{\mathrm{dof}}(\mathcal{P})=\begin{cases}\begin{aligned} &\frac{\Omega_{d-1}\ell_{\textrm{\tiny dS}}^{d-1}}{4G_{N}}\left(\frac{4\pi}{d^{2}}\right)^{d-1}\left(\tilde{\Delta}-\mathcal{P}\ell_{\textrm{\tiny dS}}\right)^{d-1}\\ &\quad\times\Bigg[1-\frac{\alpha}{\ell_{\textrm{\tiny dS}}^{2}}\,\frac{(d-1)\big(d^{2}+\mathcal{P}\ell_{\textrm{\tiny dS}}(\mathcal{P}\ell_{\textrm{\tiny dS}}+2\tilde{\Delta})\big)\left(\tilde{\Delta}-\mathcal{P}\ell_{\textrm{\tiny dS}}\right)^{2}}{3d^{2}\,(\mathcal{P}^{2}\ell_{\textrm{\tiny dS}}^{2}+d^{2})}\Bigg]\end{aligned}&\qquad\Lambda>0\ ,\\[11.99998pt] \\ \begin{aligned} &\frac{\Sigma_{k,d-1}\ell_{\textrm{\tiny AdS}}^{d-1}}{4G_{N}}\left(\frac{4\pi}{d^{2}}\right)^{d-1}\left(\mathcal{P}\ell_{\textrm{\tiny AdS}}-\Delta\right)^{d-1}\\ &\quad\times\Bigg[1+\frac{\alpha}{\ell_{\textrm{\tiny AdS}}^{2}}\,\frac{(d-1)\big(d^{2}-\mathcal{P}\ell_{\textrm{\tiny AdS}}(\mathcal{P}\ell_{\textrm{\tiny AdS}}+2\Delta)\big)\left(\mathcal{P}\ell_{\textrm{\tiny AdS}}-\Delta\right)^{2}}{3d^{2}\,(\mathcal{P}^{2}\ell_{\textrm{\tiny AdS}}^{2}-d^{2})}\Bigg]\end{aligned}&\qquad\Lambda<0\ ,\end{cases}$ (7)
where $\Omega_{d-1}=\tfrac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}$ is the area of a unit (d1)(d{-}1)-sphere, $\tilde{\Delta}\equiv\sqrt{\mathcal{P}^{2}\ell_{\textrm{\tiny dS}}^{2}+d^{2}}$ and $\Delta\equiv\sqrt{\mathcal{P}^{2}\ell_{\textrm{\tiny AdS}}^{2}-d^{2}}$. The (A)dS expressions are obtained perturbatively in the coupling $\alpha$ and they are the analytic continuation of each other, under $\ell_{\textrm{\tiny dS}}\to i\ell_{\textrm{\tiny AdS}}$. We will also provide numerical results for these expressions for finite $\alpha$.
Remarkably, it is possible to take the flat space limit, ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}\to\infty$, of the $\Lambda<0$ expression (or similarly for $\Lambda>0$) to obtain that for vanishing cosmological constant. This quantity does not receive corrections from the higher curvature terms,
@@
We also study linearized dynamics of EMGB gravity subject to ECBC in Lorentzian signature for spherical, planar and hyperbolic horizon topology. We demonstrate that the exponentially growing modes featured in Einstein gravity Anninos:2023epi ; Anninos:2024wpy ; Liu:2024ymn ; Anninos:2024xhc persist, and the GB correction does not stabilize them. Nonetheless, we observe that planar and hyperbolic black hole horizons enjoy better stability properties than the spherical ones, even within Einstein-Maxwell (EM) theory. In particular, for black branes of sufficiently large charge and for hyperbolic black holes with any value of the charge, all linearized modes are stable, except when the timelike boundary is placed very close to the horizon.
@@
One may alternatively view the action (9) as arising from the low-energy limit of a weakly coupled string theory, in which case it is possible that $\alpha\gg\ell_{p}^{2}$. In this regime, ref. Camanho:2014apa showed that tree-level three-point graviton scattering in Einstein-Gauss-Bonnet gravity alone exhibits acausal behavior. The proposed resolution is the appearance of an infinite tower of massive higher-spin states at the scale $\alpha^{1/2}$. The suggested resolution is that EMGB gravity is only consistent when embedded in a string-theoretic completion, and that one must proceed with caution, as stringy effects may become important in unexpected ways.
In the following, we present analytic (and some numerical) results for small but finite $\alpha$, treating EMGB gravity as a useful theoretical laboratory for testing ideas related to conformal boundary conditions in gravity. This perspective has proven fruitful in the past, e.g., see Myers:2010ru ; Myers:2010tj .
Secondly, even though the formulas presented in this paper will be for arbitrary dimension $d$, we would like to emphasize that it is not, even in principle, clear how these could correspond to the large-$N$ limit of some CFT when the boundary dimension $d\geq 7$. This is because Nahm’s classification Nahm:1977tg ; Minwalla:1997ka ; Cordova:2016emh prohibits the existence of superconformal algebras in $d\geq 7$, a necessary ingredient in all the known top down constructions of AdS/CFT from string theory.
@@ 2 General Framework
In the first line, we have the usual Einstein-Maxwell action with a cosmological constant. The next line is the Gauss-Bonnet (GB) term with a coupling $\alpha$, which has dimensions of length2. In four dimensions (i.e., $d=3$), the GB term is topological and so we work in $d\geq 4$ in order to have non-trivial (thermo)dynamics. [^3] One can have nontrivial thermodynamics with four bulk dimensions, since the topological GB term will contribute an additive constant to the free energy or the entropy in this case, e.g., Bunch1981 ; Sarkar:2010xp . The last line contains various boundary terms. In particular, the third term involving the Maxwell field ensures that we are working in the canonical ensemble, where the charge of the black hole is fixed – see appendix A.3 for details.
@@
In this case, the $\mathcal{K}$ boundary term is the Gibbons-Hawking-York (GHY) term York:1972sj ; Gibbons:1976ue and the $\mathcal{Q}$ term is the standard boundary term derived for Gauss-Bonnet gravity Bunch1981 ; Myers:1987yn . Other types of boundary conditions, such as Neumann or Umbilic Fournodavlos:2021eye can also be generalized for GB gravity, as explained in appendix A.
@@ 2.1 Regular Euclidean solutions
That is, with $k=\pm 1$, the above line element corresponds to the metric of a (dd–1)-sphere or (dd–1)-dimensional hyperbolic plane with unit curvature radius. With $k=0$, we have a flat metric and we have introduced a length scale $\ell$ because the $x$-coordinates are dimensionful. [^4] In the case of AdS, this $\ell$ coincides with the one appearing in the cosmological constant term of the action (9). The metric (15) is a solution of the EMGB field equations provided Boulware:1985wk ; Padmanabhan:2013xyr ; Cai:2001dz ; WILTSHIRE198636
@@
The full solution also includes the Maxwell $U(1)$ connection [^6] Recall that upon analytically continuing the charged black hole to Euclidean signature, the gauge field acquires an imaginary time component, e.g., see Gibbons:1976ue .
@@
where $\Sigma_{k,d-1}$ is the volume of maximally symmetric horizon (16). In the presence of a finite timelike boundary these will be instead related to quasi-local charges defined at the finite boundary Odak:2021axr ; Anninos:2024xhc .
In eq. (17), $f(r)$ in fact admits two branches, corresponding to choosing either the plus or minus sign in front of the square root. However, perturbing around the plus branch, one finds that the gravitons are ghosts Boulware:1985wk ; Myers:2010ru . Further, it is only the minus branch that has a smooth limit to Einstein gravity, when $\alpha\to 0$. In this limit, eq. (17) reduces to the familiar solution
@@
Before proceeding, let us comment on the GB coupling $\alpha$. In low-energy effective actions coming from string theory, the GB coupling is always positive definite Boulware:1985wk . Note also that if ${\alpha}<0$, the metric (17) has branch cut singularities, but these are always shielded by event horizons (except for small spherical black holes, where the branch cut is at the horizon) Wiltshire:1988uq . Further, with negative $\Lambda$, proper AdS vacua only appear when $4\alpha\leq\ell^{2}$, since for larger values the square root in the blackening factor (17) becomes imaginary. Holographic studies of EGB gravity Camanho:2009vw ; Buchel:2009sk yielded further constraints on $\alpha$}, by requiring consistency with the unitarity, causality and positivity of various energy fluxes in the dual CFT. These constraints, take the form Buchel:2009sk
@@
However, as noted in the introduction, the consistency of high-energy scattering processes indicates that EMGB theory is incomplete for any non-vanishing $\alpha$} Camanho:2014apa . With these caveats in mind, in the following we focus on perturbative expansions for small $\alpha$} but also take advantage of the exact solution (17) to examine results for small but finite values of the coupling.

### v2

@@ 1 Introduction
This intuition can be made precise in various top-down constructions of AdS/CFT from string theory, the prototypical example being the duality [^1] At large-$N$ and large ’t Hooft coupling $\lambda$. between a type IIB supergravity AdS black brane and $\mathcal{N}=4$ super Yang-Mills with $SU(N)$ gauge group at finite temperature Maldacena (1998); Gubser et al. (1998); Witten (1998a, b). In this example, one finds $N^{2}\propto\ell_{\textrm{\tiny AdS}}^{d-1}/G_{N}$.
From the bulk point of view, the thermodynamic behavior of black holes can be analyzed with the Euclidean gravitational path integral via the standard Gibbons-Hawking prescription Gibbons and Hawking (1977). This prescription can also accommodate situations where one is interested in (quasi-local) thermodynamics of finite spacetime regions, by for instance placing the black hole inside a finite boundary as in the classic work of York York (1986). In that particular case, spherical Dirichlet boundary conditions were imposed at the boundary.
Surprisingly enough, however, Dirichlet boundary conditions, where one fixes the profile of the induced metric along the boundary $\partial\mathcal{M}$, do not generically yield a well-posed (Initial) Boundary Value Problem (IBVP) for general relativity (GR) in (Lorentzian) Euclidean signature Anderson (2008, 2010); Anninos et al. (2023b). Furthermore, in Euclidean signature, they also fail to be elliptic, meaning that quantum perturbation theory around a classical solution is ill-defined Avramidi and Esposito (1998); Witten (2021). [^2] Note, however, that certain classes of boundary geometries exhibit better behavior under Dirichlet boundary conditions, provided the Brown–York stress tensor defines a Lorentzian metric on the boundary of the same signature as the bulk An and Anderson (2025b). An analogue statement holds in Euclidean signature Anderson (2010); Witten (2021). These include e.g., spherical boundaries in de Sitter space Coleman et al. (2022); Silverstein and Torroba (2025).
An interesting choice of boundary conditions are the conformal boundary conditions (CBC) York (1972), whereby the conformal structure of the boundary $h_{mn}]$ along with the trace of the extrinsic curvature $\mathcal{K}$ are held fixed,
@@
Conformal boundary conditions yield a well-posed and elliptic Euclidean boundary problem Anderson (2008, 2010). The Lorentzian problem is more intricate, but has better existence and uniqueness properties than the Dirichlet case. See, e.g., An and Anderson (2020, 2021, 2025a, 2025b, 2025c); Liu et al. (2025) for recent developments.
The thermodynamic properties of black holes in Einstein gravity subject to CBC have been investigated with $d=2$ and 3 in Anninos et al. (2023a, 2024); Galante et al. (2025) and the analysis was then generalized to arbitrary dimension $d$ in Banihashemi et al. (2025a, b). When discussing Euclidean thermodynamics subject to CBC, the canonical thermal ensemble is defined by a conformally invariant temperature $\tilde{\beta}$ and the trace of the extrinsic curvature $\mathcal{K}$ and hence the Euclidean problem gets modified, compared to its Dirichlet counterpart. One of the main outcomes of Anninos et al. (2023a, 2024); Galante et al. (2025); Banihashemi et al. (2025a, b) was that, in the small-$\tilde{\beta}$ limit, the conformal entropy has a structure that coincides with the expectations for the thermal effective action of a local CFT in one lower dimension – in contrast with Dirichlet boundary conditions. More specifically, the leading contribution to the black hole conformal entropy in the high-temperature expansion was found to be
@@
$\mathcal{N}_{\text{dof}}^{(\text{GR})}(\mathcal{K})=\begin{cases}\dfrac{\Sigma_{k,d-1}\ell_{\textrm{\tiny dS}}^{d-1}}{4G_{N}}\left(\dfrac{4\pi}{d^{2}}\right)^{d-1}\left(\sqrt{\mathcal{K}^{2}\ell_{\textrm{\tiny dS}}^{2}+d^{2}}-\mathcal{K}\ell_{\textrm{\tiny dS}}\right)^{d-1}&\quad\Lambda>0\ ,\\[8.61108pt] \dfrac{\Sigma_{k,d-1}\left(2\pi\right)^{d-1}}{4G_{N}\mathcal{K}^{d-1}}\ &\quad\Lambda=0\ ,\\[8.61108pt] \dfrac{\Sigma_{k,d-1}\ell_{\textrm{\tiny AdS}}^{d-1}}{4G_{N}}\left(\dfrac{4\pi}{d^{2}}\right)^{d-1}\left(\mathcal{K}\ell_{\textrm{\tiny AdS}}-\sqrt{\mathcal{K}^{2}\ell_{\textrm{\tiny AdS}}^{2}-d^{2}}\right)^{d-1}\ &\quad\Lambda<0\ .\end{cases}$ (5)
where $\Lambda=\pm\tfrac{d(d-1)}{2\ell_{\textrm{\tiny(A)dS}}^{2}}$ and $\Sigma_{k,d-1}$ is the volume of the (d1)(d-1)-dimensional maximally symmetric black hole horizon. Note that in $d=2$, and for black branes, these formulas are exact. This observation, drawing inspiration from the traditional AdS/CFT examples (1), could possibly provide hints towards a version of holography for finite portions of spacetime, even beyond $\Lambda<0$. A tree-level proposal for the dual field theoretic description of these timelike boundaries in AdS3, that reproduces eq. (5), was recently proposed in Allameh and Shaghoulian (2025b).
Given the above, a well-motivated question is to study the universality of the extensive behavior of the conformal entropy (4), when higher curvature contributions to the gravitational effective action are taken into account. In this work, we take a step towards addressing this question by considering Einstein-Maxwell-Gauss-Bonnet (EMGB) gravity. Further, we provide arguments regarding the robustness of our conclusions against further higher curvature terms of Lanczos-Lovelock (LL) type Lanczos (1938); Lovelock (1971).
@@
Studying the thermodynamics of EMGB black holes with spherical maximally symmetric horizons subject to the ECBC, we find that the extensive form of the conformal entropy in the high-temperature limit (4) persists, with (5) generalizing to
$\mathcal{N}_{\mathrm{dof}}(\mathcal{P})=\begin{cases}\begin{aligned} &\frac{\Sigma_{k,d-1}\ell_{\textrm{\tiny dS}}^{d-1}}{4G_{N}}\left(\frac{4\pi}{d^{2}}\right)^{d-1}\left(\tilde{\Delta}-\mathcal{P}\ell_{\textrm{\tiny dS}}\right)^{d-1}\\ &\quad\times\Bigg[1-\frac{\alpha}{\ell_{\textrm{\tiny dS}}^{2}}\,\frac{(d-1)\big(d^{2}+\mathcal{P}\ell_{\textrm{\tiny dS}}(\mathcal{P}\ell_{\textrm{\tiny dS}}+2\tilde{\Delta})\big)\left(\tilde{\Delta}-\mathcal{P}\ell_{\textrm{\tiny dS}}\right)^{2}}{3d^{2}\,(\mathcal{P}^{2}\ell_{\textrm{\tiny dS}}^{2}+d^{2})}\Bigg]\end{aligned}&\qquad\Lambda>0\ ,\\[11.99998pt] \\ \begin{aligned} &\frac{\Sigma_{k,d-1}\ell_{\textrm{\tiny AdS}}^{d-1}}{4G_{N}}\left(\frac{4\pi}{d^{2}}\right)^{d-1}\left(\mathcal{P}\ell_{\textrm{\tiny AdS}}-\Delta\right)^{d-1}\\ &\quad\times\Bigg[1+\frac{\alpha}{\ell_{\textrm{\tiny AdS}}^{2}}\,\frac{(d-1)\big(d^{2}-\mathcal{P}\ell_{\textrm{\tiny AdS}}(\mathcal{P}\ell_{\textrm{\tiny AdS}}+2\Delta)\big)\left(\mathcal{P}\ell_{\textrm{\tiny AdS}}-\Delta\right)^{2}}{3d^{2}\,(\mathcal{P}^{2}\ell_{\textrm{\tiny AdS}}^{2}-d^{2})}\Bigg]\end{aligned}&\qquad\Lambda<0\ ,\end{cases}$ (7)
where we defined $\tilde{\Delta}\equiv\sqrt{\mathcal{P}^{2}\ell_{\textrm{\tiny dS}}^{2}+d^{2}}$ and $\Delta\equiv\sqrt{\mathcal{P}^{2}\ell_{\textrm{\tiny AdS}}^{2}-d^{2}}$. The (A)dS expressions are obtained perturbatively in the coupling $\alpha$ and they are the analytic continuation of each other, under $\ell_{\textrm{\tiny dS}}\to i\ell_{\textrm{\tiny AdS}}$. We will also provide numerical results for these expressions for finite $\alpha$.
Remarkably, it is possible to take the flat space limit, ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}\to\infty$, of the $\Lambda<0$ expression (or similarly for $\Lambda>0$) to obtain that for vanishing cosmological constant this quantity does not receive corrections from the higher curvature terms,
@@
We also study linearized dynamics of EMGB gravity subject to ECBC in Lorentzian signature for spherical, planar and hyperbolic horizon topology. We demonstrate that the exponentially growing modes featured in Einstein gravity Anninos et al. (2023a, 2024); Liu et al. (2024); Anninos et al. (2025a, b) persist, and the GB correction does not stabilize them. Nonetheless, we observe that planar and hyperbolic black hole horizons enjoy better stability properties than the spherical ones, even within Einstein-Maxwell (EM) theory. In particular, for black branes of sufficiently large charge and for hyperbolic black holes with any value of the charge, all linearized modes are stable, except when the timelike boundary is placed very close to the horizon.
@@
One may alternatively view the action (9) as arising from the low-energy limit of a weakly coupled string theory, in which case it is possible that $\alpha\gg\ell_{p}^{2}$. In this regime, ref. Camanho et al. (2016) showed that tree-level three-point graviton scattering in Einstein-Gauss-Bonnet gravity alone exhibits acausal behavior. The proposed resolution is the appearance of an infinite tower of massive higher-spin states at the scale $\alpha^{1/2}$. The suggested resolution is that EMGB gravity is only consistent when embedded in a string-theoretic completion, and that one must proceed with caution, as stringy effects may become important in unexpected ways.
In the following, we present analytic (and some numerical) results for small but finite $\alpha$, treating EMGB gravity as a useful theoretical laboratory for testing ideas related to conformal boundary conditions in gravity. This perspective has proven fruitful in the past, e.g., see Myers and Robinson (2010); Myers and Sinha (2011).
Secondly, even though the formulas presented in this paper will be for arbitrary dimension $d$, we would like to emphasize that it is not, even in principle, clear how these could correspond to the large-$N$ limit of some CFT when the boundary dimension $d\geq 7$. This is because Nahm’s classification Nahm (1978); Minwalla (1998); Cordova et al. (2019) prohibits the existence of superconformal algebras in $d\geq 7$, a necessary ingredient in all the known top down constructions of AdS/CFT from string theory.
@@ 2 General Framework
In the first line, we have the usual Einstein-Maxwell action with a cosmological constant. The next line is the Gauss-Bonnet (GB) term with a coupling $\alpha$, which has dimensions of length2. In four dimensions (i.e., $d=3$), the GB term is topological and so we work in $d\geq 4$ in order to have non-trivial (thermo)dynamics. [^3] One can have nontrivial thermodynamics with four bulk dimensions, since the topological GB term will contribute an additive constant to the free energy or the entropy in this case, see e.g., Bunch (1981); Liko (2008); Sarkar and Wall (2011). The last line contains various boundary terms. In particular, the third term involving the Maxwell field ensures that we are working in the canonical ensemble, where the charge of the black hole is fixed – see appendix A.3 for details.
@@
In this case, the $\mathcal{K}$ boundary term is the Gibbons-Hawking-York (GHY) term York (1972); Gibbons and Hawking (1977) and the $\mathcal{Q}$ term is the standard boundary term derived for Gauss-Bonnet gravity Bunch (1981); Myers (1987a). Other types of boundary conditions, such as Neumann or Umbilic Fournodavlos and Smulevici (2023) can also be generalized for GB gravity, as explained in appendix A.
@@ 2.1 Regular Euclidean solutions
That is, with $k=\pm 1$, the above line element corresponds to the metric of a (dd–1)-sphere or (dd–1)-dimensional hyperbolic plane with unit curvature radius. With $k=0$, we have a flat metric and we have introduced a length scale $\ell$ because the $x$-coordinates are dimensionful. [^4] In the case of AdS, this $\ell$ coincides with the one appearing in the cosmological constant term of the action (9). The metric (15) is a solution of the EMGB field equations provided Boulware and Deser (1985); Padmanabhan and Kothawala (2013); Cai (2002); Wiltshire (1986)
@@
The full solution also includes the Maxwell $U(1)$ connection [^6] Recall that upon analytically continuing the charged black hole to Euclidean signature, the gauge field acquires an imaginary time component, e.g., see Gibbons and Hawking (1977).
@@
where $\Sigma_{k,d-1}$ is the volume of maximally symmetric horizon (16). In the presence of a finite timelike boundary these will be instead related to quasi-local charges defined at the finite boundary Odak and Speziale (2021); Anninos et al. (2025a).
In eq. (17), $f(r)$ in fact admits two branches, corresponding to choosing either the plus or minus sign in front of the square root. However, perturbing around the plus branch, one finds that the gravitons are ghosts Boulware and Deser (1985); Myers and Robinson (2010). Further, it is only the minus branch that has a smooth limit to Einstein gravity, when $\alpha\to 0$. In this limit, eq. (17) reduces to the familiar solution
@@
Before proceeding, let us comment on the GB coupling $\alpha$. In low-energy effective actions coming from string theory, the GB coupling is always positive definite Boulware and Deser (1985). Note also that if ${\alpha}<0$, the metric (17) has branch cut singularities, but these are always shielded by event horizons (except for small spherical black holes, where the branch cut is at the horizon) Wiltshire (1988). Further, with negative $\Lambda$, proper AdS vacua only appear when $4\alpha\leq\ell^{2}$, since for larger values the square root in the blackening factor (17) becomes imaginary. Holographic studies of EGB gravity Camanho and Edelstein (2010); Buchel et al. (2010) yielded further constraints on $\alpha$}, by requiring consistency with the unitarity, causality and positivity of various energy fluxes in the dual CFT. These constraints, take the form Buchel et al. (2010)
@@
However, as noted in the introduction, the consistency of high-energy scattering processes indicates that EMGB theory is incomplete for any non-vanishing $\alpha$} Camanho et al. (2016). With these caveats in mind, in the following we focus on perturbative expansions for small $\alpha$} but also take advantage of the exact solution (17) to examine results for small but finite values of the coupling.